# Holographic Boundary Conformal Field Theory within Horndeski Gravity

### v1

@@ I Introduction
Hence, developments in holographic transport and gravity-related quantum effects present a comprehensive overview of the intricate connections between the Anti-de-Sitter/Conformal Field Theory (AdS/CFT) correspondence and modified gravity theories, particularly Horndeski gravity Maldacena:1997re ; Witten:1998qj ; Santos:2024cwf ; Takayanagi:2011zk ; Fujita:2011fp ; Kanda:2023zse ; dosSantos:2022scy . Moreover, the extension of AdS/CFT proposed by Takayanagi, which emphasizes boundary effects and their relation to entanglement entropy Takayanagi:2011zk ; Fujita:2011fp ; Ryu:2006bv ; DosSantos:2022exb , opens promising avenues for exploring the interplay between quantum gravity and field theories Karch:2000ct ; DeWolfe:2001pq ; Bak:2003jk ; Clark:2004sb ; Cardy:2004hm ; Azeyanagi:2007qj ; Setare:2008hm ; Saridakis:2017rdo ; Basilakos:2023seo . The discussion of Hawking-Page phase transitions and corrections to boundary entropy underscores the richness of the holographic framework in addressing complex phenomena. By describing gravity duals within this context enlightens the discussion on the relation between quantum information theory and gravitational physics, particularly in scenarios where traditional quantum gravity approaches may be inadequate.
@@
The recent discussions on doubly holographic descriptions and the island formula reveal the interplay between holography, entanglement, and gravity. The use of BCFTs to understand these concepts is a significant advancement, especially in the context of understanding how entanglement entropy behaves in gravitational settings Almheiri:2019hni ; Rozali:2019day ; Chen:2020uac ; Chen:2020hmv ; Deng:2020ent ; Suzuki:2022xwv ; Geng:2021iyq ; Geng:2020fxl . The framework we are outlining, where a CFT coupled to semiclassical gravity provides a dual description of a (d+1)-dimensional theory, is a robust platform for exploring entanglement phenomena. Additionally, the Ryu-Takayanagi (RT) prescription is central to this exploration Ryu:2006ef ; Ryu:2006bv , enabling the calculation of holographic entanglement entropy in terms of geometric quantities in the gravitational dual Suzuki:2022xwv .
@@
The analysis of the interplay between a CFTd coupled to AdSd gravity on the brane, particularly in the context of transparent boundary conditions, offers information on the dynamics of such systems Basu:2023jtf ; Hartman:2013qma ; Geng:2021mic ; Geng:2023qwm ; Geng:2021hlu ; Geng:2023zhq ; Geng:2022dua . These boundary conditions serve as a crucial interface between the brane dynamics and the non-gravitational bath, leading to significant physical phenomena, such as the effective mass acquisition of the graviton due to energy exchange Karch:2000ct ; Brito:2018pwe ; Santos:2023eqp . The role of transparent boundary conditions in facilitating this mass acquisition is important, as highlighted in works like Geng:2021mic . This mechanism illustrates how the gravitational theory on the brane can influence bulk dynamics, thereby affecting the properties of gravitational excitations.
@@
This work is summarized as follows. In Sec. II we present the doubly holographic black string and we extract the conditions that this solution satisfies to become a solution of the Horndeski gravity. In Sec. III we calculate the density functional to provide the entanglement thermodynamics in Horndeski gravity, and we extract the entanglement entropy through spatial embedding formalism and Field Theory computation, while in Sec. IV we present the Page curve behavior. Finally, Sec. V provides our conclusions.
II Horndeski gravity on doubly holographic black string
@@
The exploration of the implications of Dirichlet and Neumann boundary conditions is important for understanding the construction of of RT surfaces, and can lead to different physical interpretations regarding the flow of information and the stability of entangled states in the boundary theory. Nevertheless, the different angles of the brane, particularly below the critical angle (see Fig. 4), introduces more complications. In particular, the critical angle can mark a transition in the behavior of the RT surfaces, possibly affecting their stability and the nature of the entanglement islands, and surfaces that connect to the brane at angles below this critical threshold might exhibit different entanglement properties than those that do not.
@@
III Entanglement entropy and Hatman-Maldacena surface
@@
In this section we present the entanglement entropy to the subregion $\mathcal{R}$, $\bar{\mathcal{R}}$ and defect, which leads to the island on the brane. In general, it is possible to find extremal surfaces $\Omega$ that satisfy the holonomy constraint like $\partial\,\Omega=\partial\mathcal{\mathcal{R}}\cup\partial\bar{\mathcal{R}}$. In scenarios involving defects or branes, the presence of an island can complicate the calculation of entanglement entropy. The interaction between the two islands is essential for ensuring that the entanglement entropy adheres to unitarity. As the system evolves, the total entropy reflects information conservation. The dominance of an island surface over an Hatman-Maldacena (HM) Hartman:2013qma surface at an initial time indicates that the entanglement dynamics can shift, revealing the non-trivial evolution of entanglement in these scenarios. Page time is particularly relevant here, as it marks the moment at which the entanglement entropy transit from being dominated by the HM surface to the island surface. The dependence of Page time on parameters such as $u_{b}$, $\alpha$, $\gamma$, and the sizes $r_{L}$ and $r_{R}$, indicates the richness of the model. Each parameter plays a role in determining the degrees of freedom in the defect and the size of the regions of interest Takayanagi:2011zk . Finally, the observation that Page time is approximately proportional to the difference in area between the RT surfaces at $t=0$, reveals a geometric aspect of the entanglement dynamics, and this area difference reflects the competing contributions to the entanglement entropy and significantly determines how the entropy evolves.
The Ryu-Takayanagi formula Ryu:2006ef is
$S=\frac{\mathcal{A}(\partial\mathcal{R})}{4G}+\frac{\mathcal{A}({\partial\bar{\mathcal{R}}})}{4G},$ (20)
where $\mathcal{A}(\partial\mathcal{R})$ and $\mathcal{A}({\partial\bar{\mathcal{R}}})$ are the areas given by

### v2

@@ 1 Introduction
Hence, developments in holographic transport and gravity-related quantum effects present a comprehensive overview of the intricate connections between the Anti-de-Sitter/Conformal Field Theory (AdS/CFT) correspondence and modified gravity theories, particularly Horndeski gravity Maldacena:1997re ; Witten:1998qj ; Caceres:2023gfa ; Santos:2024cwf ; Takayanagi:2011zk ; Fujita:2011fp ; Kanda:2023zse ; dosSantos:2022scy . Moreover, the extension of AdS/CFT proposed by Takayanagi, which emphasizes boundary effects and their relation to entanglement entropy Takayanagi:2011zk ; Fujita:2011fp ; Ryu:2006bv ; DosSantos:2022exb ; Caceres:2017lbr ; Feng:2015oea ; Wald:1993nt ; Iyer:1994ys , opens promising avenues for exploring the interplay between quantum gravity and field theories Karch:2000ct ; DeWolfe:2001pq ; Bak:2003jk ; Clark:2004sb ; Cardy:2004hm ; Tonni:2010pv ; Azeyanagi:2007qj ; Setare:2008hm ; Saridakis:2017rdo ; Basilakos:2023seo . The discussion of Hawking-Page phase transitions and corrections to boundary entropy underscores the richness of the holographic framework in addressing complex phenomena. By describing gravity duals within this context enlightens the discussion on the relation between quantum information theory and gravitational physics, particularly in scenarios where traditional quantum gravity approaches may be inadequate.
@@
The recent discussions on doubly holographic descriptions and the island formula reveal the interplay between holography, entanglement, and gravity. The use of BCFTs to understand these concepts is a significant advancement, especially in the context of understanding how entanglement entropy behaves in gravitational settings Almheiri:2019hni ; Rozali:2019day ; Chen:2020uac ; Chen:2020hmv ; Deng:2020ent ; Suzuki:2022xwv ; Geng:2021iyq ; Geng:2024xpj ; Geng:2020fxl . The framework we are outlining, where a CFT coupled to semiclassical gravity provides a dual description of a (d+1)-dimensional theory, is a robust platform for exploring entanglement phenomena. Additionally, the Ryu-Takayanagi (RT) prescription is central to this exploration Ryu:2006ef ; Ryu:2006bv , enabling the calculation of holographic entanglement entropy in terms of geometric quantities in the gravitational dual Suzuki:2022xwv .
@@
The analysis of the interplay between a CFTd coupled to AdSd gravity on the brane, particularly in the context of transparent boundary conditions, offers information on the dynamics of such systems Basu:2023jtf ; Hartman:2013qma ; Geng:2024xpj ; Geng:2021mic ; Geng:2023qwm ; Geng:2021hlu ; Geng:2023zhq ; Geng:2022dua . These boundary conditions serve as a crucial interface between the brane dynamics and the non-gravitational bath, leading to significant physical phenomena, such as the effective mass acquisition of the graviton due to energy exchange Karch:2000ct ; Brito:2018pwe ; Santos:2023eqp . The role of transparent boundary conditions in facilitating this mass acquisition is important, as highlighted in works like Geng:2021mic . This mechanism illustrates how the gravitational theory on the brane can influence bulk dynamics, thereby affecting the properties of gravitational excitations.
@@
This work is summarized as follows. In Sec. 2 we present the doubly holographic black string and we extract the conditions that this solution satisfies to become a solution of the Horndeski gravity. In Sec. 3 we calculate the density functional to provide the entanglement thermodynamics in Horndeski gravity, and we extract the entanglement entropy through spatial embedding formalism and Field Theory computation, while in Sec. 4 we present the Page curve behavior. Finally, Sec. 5 provides our conclusions.
2 Horndeski gravity on doubly holographic black string
@@
The exploration of the implications of Dirichlet and Neumann boundary conditions is important for understanding the construction of RT surfaces, and can lead to different physical interpretations regarding the flow of information and the stability of entangled states in the boundary theory. Nevertheless, the different angles of the brane, particularly below the critical angle (see Fig. 4), introduces more complications. In particular, the critical angle can mark a transition in the behavior of the RT surfaces, possibly affecting their stability and the nature of the entanglement islands, and surfaces that connect to the brane at angles below this critical threshold might exhibit different entanglement properties than those that do not.
@@
3 Entanglement entropy and Hatman-Maldacena surface
@@
In this section we present the entanglement entropy to the subregion $\mathcal{R}$, $\bar{\mathcal{R}}$ and defect, which leads to the island on the brane. In general, it is possible to find extremal surfaces $\Omega$ that satisfy the holonomy constraint like $\partial\,\Omega=\partial\mathcal{\mathcal{R}}\cup\partial\bar{\mathcal{R}}$. In scenarios involving defects or branes, the presence of an island can complicate the calculation of entanglement entropy. The interaction between the two islands is essential for ensuring that the entanglement entropy adheres to unitarity. As the system evolves, the total entropy reflects information conservation. The dominance of an island surface over an Hatman-Maldacena (HM) Hartman:2013qma surface at an initial time indicates that the entanglement dynamics can shift, revealing the non-trivial evolution of entanglement in these scenarios. Page time is particularly relevant here, as it marks the moment when the entanglement entropy transit from being dominated by the HM surface to the island surface. The dependence of Page time on parameters such as $u_{b}$, $\alpha$, $\gamma$, and the sizes $r_{L}$ and $r_{R}$, indicates the richness of the model. Each parameter plays a role in determining the degrees of freedom in the defect and the size of the regions of interest Takayanagi:2011zk . Finally, the observation that Page time is approximately proportional to the difference in area between the RT surfaces at $t=0$, reveals a geometric aspect of the entanglement dynamics, and this area difference reflects the competing contributions to the entanglement entropy and significantly determines how the entropy evolves.
Before presenting the entanglement entropy and Hatman-Maldacena surface in Horndeski’s scenario. We will present entanglement entropy calculation in the Karch-Randall braneworld in Sec. 3.1, it emerges naturally as presented by Geng:2024xpj . Based on this scenario, we can extend the calculation of entanglement entropy to the Horndeski scenario Sec. 3.2.
3.1 Entanglement entropy calculation in the Karch-Randall braneworld
The Karch-Randall braneworld scenario provided a natural setting to study Hawking radiation from a black hole using holographic tools Randall:1999vf ; Karch:2000ct ; Brito:2018pwe ; Santos:2023eqp . However, quantum effects were an open problem until recently, leaving a gap that needs to be filled. This gap was filled in the work Geng:2024xpj where the ambient space is (2+1)-dimensional, for which explicit calculations can be made in each configuration description.
As discussed and shown by Geng:2024xpj , the Karch-Randall braneworld model involves embedding a subcritical brane within an ambient asymptotically AdSd+1 spacetime, denoted as $\mathcal{N}_{d+1}$. In this framework, the ambient spacetime is governed by classical Einstein gravity, while the brane is a codimension-one hypersurface, $\mathcal{N}_{d}$. The action governing this system is given by
$S=-\frac{1}{16\pi G_{d+1}}\int_{\mathcal{N}_{d+1}}d^{d+1}x\sqrt{-g}(R-2\Lambda)-\frac{1}{8\pi G_{d+1}}\int_{\mathcal{N}_{d}}d^{d}x\sqrt{-h}(K-T)\,,$ (20)
The boundary conditions for the bulk metric in the Karch-Randall braneworld model are of the Neumann type near the brane and the Dirichlet type near the remaining asymptotic boundary Geng:2023qwm ; Geng:2021mic . This implies nonzero metric fluctuations near the brane, while fluctuations vanish near the asymptotic boundary. In this framework, the brane represents a gravitational AdSd+1 spacetime, whereas the asymptotic boundary acts as a $d$-dimensional nongravitational bath Geng:2021iyq ; Geng:2020fxl .
Two sets of equations of motion govern the system:
  • Einstein’s equations, which ensure the vanishing of the bulk variation of the action (20), and the brane embedding equations, vanishing to the brane for the variation of the same action;
  • Einstein’s equations determine the background geometry of the ambient spacetime. In contrast, the brane embedding equations dictate how the brane is embedded within this spacetime, thereby defining the induced geometry on the brane Santos:2021orr ; Geng:2024xpj .
For the bulk description is asymptotically AdS3 with Karch-Randall braneworld Karch:2000ct , this description is associated to BCFT2; we want to calculate the subregion entanglement entropy SR of the thermal field dual state Geng:2021mic . To study the thermal field double states $|TFD\rangle$, one intrudes two Karch-Randall branes (see Fig. 6) into the spacetime of the ambient BTZ black hole; the BCFT2 are in this double state. In this boundary description, the two BCFT2 live in strips with boundary conditions corresponding to Cardy states Cardy:2004hm .
Figure 6: Configuration with two branes in the BTZ background where the gray regions represent cut off by the branes.
In the setup illustrated in Fig. 7 Geng:2024xpj , we consider a model involving two eternal black holes in (1+1)-dimensions, which are coupled through a thermal bath (see Fig. 7-a). This configuration is analyzed within the framework of Boundary Conformal Field Theory (BCFT), where the boundaries are characterized by specific boundary conditions associated with Cardy states (refer to Fig. 7-b) Ryu:2006bv ; Cardy:2004hm ; Geng:2024xpj . This setup provides a novel perspective on the interaction dynamics between black holes and their environments, offering insights into the underlying quantum gravitational processes.
Figure 7: The figure shows in (a) Penrose diagram of the intermediate description for two branes at the bottom of the BTZ black hole; we have two ($1+1$)-dimensional black holes, which are coupled to each other through strip-shaped baths (the green shaded region). We identify the two outer edges in red. It can be seen that the black hole singularities are orbifold singularities inherited from the BTZ black holes. In part of frame (b), we have a description of the boundary of the configuration with two branes at the bottom of the BTZ black hole.
Through the holographic dictionary of the Karch-Randall braneworld, one can perform a mapping of the replicated boundary partition function into a Euclidean gravitational path integral as follows:
$Z_{\text{BCFT}}[\mathcal{N}^{(n)}]=\int_{\partial\mathcal{N}^{(n)}_{d+1}=\mathcal{N}^{(n)}}D[g]e^{-S}=Z_{\text{grav}}[\mathcal{N}^{(n)}]\,,$ (21)
[…]
$=$ $\lim_{n\rightarrow 1}\frac{1}{1-n}\frac{Z_{\text{grav}}[\mathcal{N}^{(n)}]}{Z_{\text{grav}}[\mathcal{N}^{(0)}]^{n}}\,,$ (23)
[…]
Figure 8: The diagram illustrates the Euclidean path integral formulation with an initially prepared ground state, denoted as $|0\rangle$. Within the dashed interval, we identify the subregion $\mathcal{R}$. The reduced density matrix operator for this subregion, $\hat{\rho}_{\mathcal{R}}$, is derived from the full density matrix $\hat{\rho}=|0\rangle\langle 0|$ by integrating over all possible field configurations outside of $\mathcal{R}$. This approach allows us to isolate the contributions specific to the subregion, providing insights into the entanglement properties and quantum correlations present in the system.
Through the half Euclidean plane with the Cardy boundary-$\mathcal{N}$ Ryu:2006bv ; Cardy:2004hm ; Geng:2024xpj , the trace of the $n-th$ power of the normalized reduced density matrix operator is given by
$tr\hat{\rho}_{\bar{R}}^{n}=\frac{Z_{BCFT}[\mathcal{N}^{(n)}]}{Z^{n}_{BCFT}[\mathcal{N}]}\,.$ (25)
This expression is the key point to prove the Ryu-Takayanagi Ryu:2006bv conjecture in the AdS/BCFT correspondence scenario using gravitational path integral. Thus, with this, we can conclude that
$S_{R}$ $=$ $\lim_{n\rightarrow 1}\frac{1}{1-n}\frac{e^{-nS[g^{BTZ}]+\frac{A(\gamma)}{4G_{d+1}}\frac{1-n}{n}]}}{e^{-nS[g^{BTZ}]}}\,,$ (26)
$=$ $\frac{A(\gamma)}{4G_{d+1}}\,.$ (27)
where $g^{EBTZ}$ is the Euclidean BTZ metric. The phases of the Karch-Randall brane within the manifold $\mathcal{N}_{d+1}^{(n)}$ Geng:2021mic ; Geng:2022dua .
With this prescription, we can identify two distinct phases:
  • First phase: characterized by $n$-disconnected components and another by a single component that connects the Cardy boundaries of the boundary replica manifold $\mathcal{N}_{d+1}^{(n)}$ Cardy:2004hm . In this phase, the conical singularity $\gamma$ remains in the bulk, linking the two $\partial\mathcal{R}$ boundaries;
  • Second phase: $\gamma$ extends to the brane, with $\mathbb{Z}_{n}$ fixed points present, resulting in two disconnected components Tonni:2010pv that connect the boundaries of $\partial\mathcal{R}$ to the brane Geng:2023qwm ; Geng:2021hlu .
This dual-phase behavior provides new insights into the geometric and topological properties of the Karch-Randall brane, offering potential implications for understanding brane dynamics in higher-dimensional spaces Geng:2022dua . However, we have that
$S(R)=\min(\frac{A_{c}}{4G_{d+1}},\frac{2A_{dc}}{4G_{d+1}})\,,$ (28)
This is the Ryu-Takayanagi conjecture Ryu:2006bv . Note that $\gamma_{dc}$ must also have its area minimized over its possible endpoints on the brane; this corroborates our saddle point approximation. This discussion shows that the Ryu-Takayanagi conjecture is proven; one can see that, indeed, the emergence of the disconnected Ryu-Takayanagi surface is due to the connected phase of the Karch-Randall brane in the replica path integral Basu:2023jtf .
3.2 Derivation of the functional entanglement entropy for Horndeski gravity
We must derive the functional area for Horndeski gravity’s entanglement entropy to study the holographic entanglement entropy. Considering the Horndeski Lagrangian ${\cal L}_{H}$, and following the technique detailed in previous works of DosSantos:2022exb ; Caceres:2017lbr ; Feng:2015oea ; Wald:1993nt ; Iyer:1994ys , which consider:
  • The Lagrangian-${\cal L}_{H}$ of Horndeski theory does not include quadratic or higher-order curvature terms;
  • Incorporates a Wald-type term Feng:2015oea , which is important in developing holographic entanglement entropy in arbitrary gravity theories.
This approach facilitates the application of the replica trick, a method essential for deriving holographic entanglement entropy. Through the microscopic definition of entanglement entropy Ryu:2006ef ; Caceres:2017lbr :
$S_{EE}=-Tr(\rho\log(\rho))$ (29)
Here, $\rho$ is the reduced density matrix of a subregion of a time slice of the boundary. Here, the idea is to find the entanglement entropy in the subregion $\mathcal{R}$, $\bar{\mathcal{R}}$. For this, we apply the replica trick through the $n^{th}$ Rényi entropy, defined by:
$S_{n}=-\frac{1}{n-1}\log{\mathrm{Tr}{(\rho^{n})}}.$ (30)
Here the entanglement entropy is the analytical continuation of $S_{n}$ as $n\rightarrow 1$: $S_{EE}=\lim_{n\rightarrow 1}S_{n}$. From the overview point of path integral, $\rho$, can be express $S_{n}$ in terms of the partition function $Z_{n}$ considering an appropriate n-sheeted Riemann surface $M_{n}$ as:
$S_{n}=-\frac{1}{n-1}\left(\log{Z_{n}}-n\log{Z_{1}}\right),$ (31)
in the above equation $Z_{1}$ is the original partition function. Besides, using the basic holographic relation $Z_{CFT}=e^{-S_{bulk}}$ between the field theory partition function $Z_{CFT}$ and the bulk action for an appropriate bulk geometry. With this and through the prescription of Caceres:2017lbr ; Feng:2015oea , we can derive the area functional for Entanglement Entropy for Horndeski gravity ($S_{EEH}$):
$S_{EEH}=\int_{\mathcal{N}}{d^{d}x\sqrt{\eta}\frac{\partial{\cal L}_{H}}{\partial R_{\mu\nu\rho\sigma}}\epsilon_{\mu\nu}\epsilon_{\rho\sigma}},$ (32)
where $\epsilon_{\mu\nu}$ and $\epsilon_{\rho\sigma}$ denotes the bi-normal to $\mathcal{N}$ Feng:2015oea ; Wald:1993nt ; Iyer:1994ys . We can derive Horndeski Lagrangian with respect to the Riemann tensor as:
$\frac{\partial{\cal L}_{H}}{\partial R_{\mu\nu\rho\sigma}}$ $=$ $-\frac{\kappa}{32\pi}(g^{\mu\rho}g^{\nu\sigma}-g^{\nu\rho}g^{\mu\sigma})-\frac{\gamma}{128\pi}[g^{\mu\rho}\phi^{,\nu}\phi^{,\sigma}-g^{\nu\rho}\phi^{,\mu}\phi^{,\sigma}+g^{\nu\sigma}\phi^{,\mu}\phi^{,\rho}-g^{\mu\sigma}\phi^{,\nu}\phi^{,\rho}$ (33)
-- $(g^{\mu\rho}g^{\nu\sigma}-g^{\nu\rho}g^{\mu\sigma})\phi^{,\lambda}\phi_{,\lambda}]$
Here in our work $\phi^{,\alpha}=g^{\alpha\beta}\phi_{,\beta}$. For the metric in complex coordinates like:
$ds^{2}=dzd\bar{z}+\eta_{ij}dy^{i}dy^{j}$ (34)
Upon substituting the metric components into (33), we have
$\frac{\partial{\cal L}_{H}}{\partial R_{z\bar{z}z\bar{z}}}=\frac{\kappa}{8\pi}-\frac{\gamma}{32\pi}\left(\phi^{,\lambda}\phi_{,\lambda}-\phi^{,z}\phi^{\bar{,z}}\right)$ (35)
Upon further expanding $\phi^{,\lambda}\phi_{,\lambda}=\phi^{,z}\phi^{,\bar{z}}+\eta_{ij}\phi^{,i}\phi^{,j}$, this further simplifies to:
$\frac{\partial{\cal L}_{H}}{\partial R_{z\bar{z}z\bar{z}}}=\frac{\kappa}{8\pi}-\frac{\gamma}{32\pi}\eta_{ij}\phi^{,i}\phi^{,j}$ (36)
Thus, we have that the functional for holographic entanglement entropy then read Caceres:2017lbr ; Feng:2015oea :
$S_{EEH}=\int{d^{d}x\sqrt{\eta}\frac{\partial{\cal L}_{H}}{\partial R_{z\bar{z}z\bar{z}}}}=\frac{\kappa}{4}\int{d^{d}x\sqrt{\eta}\left[1-\frac{\gamma}{4\kappa}\eta_{ij}\phi^{,i}\phi^{,j}\right]}$ (37)
We can note that comparing our result with the Ryu-Takayanagi formula Ryu:2006ef , the correction for the area is due to the term $\gamma\,G_{\mu\nu}\nabla^{\mu}\phi\nabla^{\nu}\phi$ where for $\gamma=0$ we recover the usual Einstein case. Considering the metric (1) and (37), we can derive the area functional $\mathcal{A}(\partial\mathcal{\mathcal{R}}\cup\partial\bar{\mathcal{R}})$ to the subregion $\mathcal{R}$ and $\bar{\mathcal{R}}$ for Horndeski gravity, which is given by
$S_{EEH}=\frac{\mathcal{A}}{4G}$ (38)